### v2

# JugglePAC: a Pipelined Accumulation Circuit

Abstract
Reducing a set of numbers to a single value is a fundamental operation in applications such as signal processing, data compression, scientific computing, and neural networks. Accumulation, which involves summing a dataset to obtain a single result, is crucial for these tasks. Due to hardware constraints, large vectors or matrices often cannot be fully stored in memory and must be read sequentially, one item per clock cycle. For high-speed inputs, such as rapidly arriving floating-point numbers, pipelined adders are necessary to maintain performance. However, pipelining introduces multiple intermediate sums and requires delays between back-to-back datasets unless their processing is overlapped. In this paper, we present JugglePAC, a novel accumulation circuit designed to address these challenges. JugglePAC operates quickly, is area-efficient, and features a fully pipelined design. It effectively manages back-to-back variable-length datasets while consistently producing results in the correct input order. Compared to the state-of-the-art, JugglePAC achieves higher throughput and reduces area complexity, offering significant improvements in performance and efficiency.
Index Terms:
Fully pipelined reduction circuits, floating-point number accumulation, field-programmable gate arrays, computer arithmetic.
I Introduction
In the realm of modern computing, the ability to efficiently reduce a set of numbers to a single value is fundamental to a wide variety of computational applications. This process, known as reduction, is crucial for tasks ranging from signal processing [1, 2], data compression [3, 4], scientific computing [5, 6], and neural networks [7, 8]. Among the various types of reduction operations, accumulation, which involves summing a dataset to produce a single result, is one of the most essential. As data complexity and scale increase, high-performance accumulation methods that handle large datasets efficiently become more essential.
Accumulation operations can be applied to both floating-point and integer data. For integer data, the use of a 3:2 compressor simplifies the process by reducing latency, making integer accumulation straightforward. However, floating-point accumulation presents more complex challenges, necessitating pipelined adders to manage high data input rates. Pipelining, while essential for maintaining throughput in high-speed applications, introduces design complexities such as managing multiple intermediate sums and potential delays with consecutive datasets.
TABLE I: Accumulation schedule for SimplePAC versus JugglePAC.
SimplePAC        JugglePAC
Input Adder        Input Adder
in1 in2 out        in1 in2 out
$a_{0}$ $a_{0}$ 0        $a_{0}$
$a_{1}$ $a_{1}$ 0        $a_{1}$ $a_{0}$ $a_{1}$
$a_{2}$ $a_{2}$ 0        $a_{2}$
$a_{3}$ $a_{0}$ $a_{3}$ $a_{0}$        $a_{3}$ $a_{2}$ $a_{3}$
$a_{4}$ $a_{1}$ $a_{4}$ $a_{1}$        $a_{4}$ $a_{0,1}$
$a_{5}$ $a_{2}$ $a_{5}$ $a_{2}$        $a_{5}$ $a_{4}$ $a_{5}$
Stall $\mathbf{a_{0,3}}$        $b_{0}$ $a_{0,1}$ $a_{2,3}$ $a_{2,3}$
$a_{0,3}$ $a_{1,4}$ $\mathbf{a_{1,4}}$        $b_{1}$ $b_{0}$ $b_{1}$
$\mathbf{a_{2,5}}$        $b_{2}$ $a_{0,1}$ $a_{2,3}$ $a_{4,5}$
       $b_{3}$ $b_{2}$ $b_{3}$ $a_{0:3}$
$a_{0,1,3,4}$ $a_{2,5}$ $a_{0,1,3,4}$        $b_{4}$ $a_{4,5}$ $a_{0:3}$ $b_{0,1}$
$b_{0}$ $b_{0}$ 0        $b_{5}$ $b_{4}$ $b_{5}$
$b_{1}$ $b_{1}$ 0        $b_{6}$ $a_{0:3}$ $a_{4,5}$ $b_{2,3}$
$b_{2}$ $b_{2}$ 0 $\mathbf{a_{0:5}}$        $b_{7}$ $b_{6}$ $b_{7}$ $\mathbf{a_{0:5}}$
Table I depicts an example of a floating-point adder-based accumulator with a pipeline latency of 3 clock cycles, while new input values are fed in every cycle. Using a simple accumulation schedule (SimplePAC), we end up with 3 subsums, 3 cycles after all data inputs are fed to the adder. The subsums are shown as $a_{0,3}$, $a_{1,4}$, and $a_{2,5}$ in bold in Table I. The simplest approach to accumulation would be not to allow a new dataset (called $b_{0:5}$) until the last addition of the current dataset is fed into the adder pipeline (i.e., introduce stalls between consecutive datasets). However, SimplePAC does not accept back-to-back datasets. Using a more intuitive schedule, we can alternate between additions from different sets to maintain a fully pipelined accumulation circuit. Such a schedule is shown in Table I using the JugglePAC schedule, which starts accumulating the data from the consecutive dataset without requiring any stalls.
This work presents JugglePAC, a novel fully pipelined accumulation circuit that optimizes both area and timing using a single floating-point adder. We implement JugglePAC and evaluate its performance across multiple FPGAs, benchmarking it against state-of-the-art solutions. The major contributions of this work are:
  1. 1.
    We propose a fully pipelined accumulation circuit, namely JugglePAC, which efficiently handles high-speed, back-to-back variable-length datasets.
  2. 2.
    JugglePAC demonstrates improvements over existing solutions by achieving higher frequencies and reducing area complexity.
  3. 3.
    Our work introduces a dynamic scheduling mechanism for variable-length datasets, addressing the challenges associated with inefficient control logic in existing designs.
  4. 4.
    We implement JugglePAC on two different target FPGAs, specifically the Xilinx XC2VP30 and XC5VLX110T, showing consistent improvements in both area and timing over the state-of-the-art.
II Related Work
Recent work on floating-point accumulation circuits has focused on optimizing area, performance, and complexity. Early designs, such as Luo and Martonosi [9], used carry-save arithmetic and delayed adders for high-performance accumulation but were not fully pipelined, leading to performance bottlenecks due to the required stalls.
Vangal [10] improved on this by introducing a pipelined structure for single-precision floating-point multiply-accumulate operations, yet managing the pipeline complexity remained a challenge. He et al. [11] proposed a group alignment algorithm to improve accuracy but struggled with scalability and efficiency for variable-length datasets.
Nagar and Bakos [12] developed a double-precision accumulator with a coalescing reduction circuit, reducing complexity but limited by its FPGA-specific design. Zhou et al. [13] introduced several designs, including the Fully Compacted Binary Tree (FCBT) and Dual Striped Adder (DSA), which managed multiple input sets but faced issues with buffer management and clock speed.
Huang and Andrews [14] proposed modular, fully pipelined architectures capable of handling arbitrary dataset sizes, though their designs faced challenges with underutilized pipelined adders and extensive buffering.
Recent designs [15, 16, 17, 18, 19] aimed at balancing area and timing performance. [19] notably improved the area-timing product but required multiple BRAMs, impacting area efficiency.
In contrast, JugglePAC offers a novel approach with a fully pipelined architecture that simplifies control logic and handles variable-length datasets efficiently. Its dynamic scheduling mechanism ensures high frequency and low area complexity, outperforming previous designs in scalability and adaptability across different FPGA architectures.
III JugglePAC
JugglePAC is a novel floating-point accumulation circuit designed to optimize performance and area. This section describes the microarchitecture of JugglePAC and its inter-dataset behavior as well as key challenges such as the ability to mix variable-length datasets.
III-A JugglePAC Microarchitecture
The JugglePAC architecture is built around a floating-point adder, which plays a crucial role in the accumulation process. The fundamental concept of JugglePAC involves scheduling the additions of serially arriving data to achieve efficient accumulation. Given that data typically arrives serially, the adder operates with a throughput of 1, performing additions every 2 cycles. This setup ensures that the adder is utilized 50% of the time, while producing results at least every 2 cycles. Consequently, a single adder is sufficient to maintain pace with the input rate when additions are scheduled effectively. JugglePAC employs a state machine with two distinct states to manage the addition process. In the first state, the inputs are directly added. In the second state, the design handles the addition of any available pair of subsums, as illustrated in Fig. 1. The state machine iterates between these two states every cycle, with an exception for datasets of odd length, where the state machine remains in state 1 for an additional cycle.
To enhance performance for high-throughput applications, JugglePAC is designed to handle back-to-back inputs, thereby avoiding data pile-ups. This approach introduces additional complexity as it necessitates the simultaneous processing of subsums from different datasets. To manage this, each subsum is assigned a unique label, an integer that increments with each new dataset (color-coded with green in Fig. 1). This labeling system is maintained using a shift register with a latency equal to that of the adder as depicted in Fig. 1 and color-coded with purple. The Matching Shift Register block matches subsums from the same dataset and distinguishes between those from different datasets. An additional pipeline stage is added before the inputs of the adder to increase the throughput, which is represented by dashes in the figure.
Scheduling additions of subsums requires efficient control logic. The Pair Identifier (PI) block, shown in Fig. 1 and color-coded with yellow, is responsible for managing this process. The PI receives the results from the adder along with their labels and schedules the additions accordingly. It uses a register for each label, storing incoming subsums and identifying pairs for addition. When a pair is identified, the PI schedules the addition and clears the corresponding register. This control logic allows JugglePAC to juggle between additions from different datasets while minimizing area requirements. The number of registers in the PI depends on the label size, with L representing the maximum number of labels. To handle situations where the adder may not always be available for scheduled additions, JugglePAC employs a FIFO buffer to temporarily store ready-to-add subsum pairs along with their labels. The FIFO is read every 2 cycles, with a maximum depth of $\lceil\log(p)\rceil$, where $p$ is the adder’s latency. The FIFO’s design, which utilizes registers, ensures efficient data management.
The state machine within JugglePAC manages the accumulation process by maintaining a low area and timing complexity. It alternates between adding serial input data and processing data from the FIFO. A sample scheduling scenario using the JugglePAC approach is illustrated in Table I, demonstrating how the accumulation of dataset $b_{0:N}$ begins before the results of dataset $a_{0:5}$ are completed. The simple control logic for the state machine and pair identification allows JugglePAC to maintain a low area complexity and critical path, outperforming the state-of-the-art.
For output identification, JugglePAC uses a counter to track the number of additions performed. The counter increments with each addition from state 1 and decrements with each addition from state 0. The system skips incrementing on the first addition of state 1, ensuring the counter returns to zero after each operation. The counter reaches a maximum value of $\lceil\frac{p+2}{2}\rceil$, where $p$ is the adder’s latency, and returns to zero upon completing the final addition. Separate counters are used for each simultaneous accumulation, with $L$ representing the maximum number of counters. The output identifier is represented in Fig. 1 and color-coded with red.
III-B Inter-Dataset Behavior
The JugglePAC architecture allows for the parallel processing of datasets, with the label size determining the maximum number of datasets that can be handled simultaneously. A smaller label size reduces area complexity but introduces a minimum dataset length requirement. When the number of datasets exceeds $L$, the circuit may mix accumulations from different datasets, leading to incorrect results. Therefore, the minimum dataset length is identified through comprehensive testing with variable-length datasets.
Increasing the label size enhances the circuit’s capability to handle more datasets but also increases the logical resources required for the PI and output identification modules. This tradeoff between area complexity and minimum dataset length is summarized in Section IV, which shows the area complexity increasing with larger label sizes.
JugglePAC’s dynamic accumulation approach results in latency that depends on the current and previous dataset lengths. For label sizes less than 3, the circuit maintains consistent latency, ensuring that the results are produced in input order. However, with larger label sizes, the minimum dataset length can lead to results being output in an order that deviates from the input sequence. Additional control logic can be used to reorder results based on the label or to output the label itself, allowing the system to identify the dataset to which the result belongs. Setting the minimum dataset length to 19 ensures that JugglePAC consistently produces ordered results. This design achieves a balance between high frequency and low area complexity, surpassing existing state-of-the-art solutions.
Figure 1: JugglePAC microarchitecture features a Floating-point Adder and key components including a state machine for managing additions, a Matching Shift Register for labeling subsums, and a Pair Identifier block for scheduling. A FIFO buffer ensures efficient data handling and prevents pile-ups, while an Output Identifier tracks the number of additions performed. This design optimizes throughput and minimizes area complexity.
IV Implementation Results
We evaluated the JugglePAC floating-point accumulation circuit against existing designs, focusing on area complexity, latency, throughput, and frequency. Table II summarizes these metrics for JugglePAC and other designs.
JugglePAC demonstrates a significant reduction in area complexity, using fewer slices than designs like MFPA, AeMFPA, and FAAC [14], with up to 71% less slice usage. Additionally, JugglePAC operates without BRAMs, unlike FCBT and DSA [13], contributing to its lower area complexity.
In terms of latency, JugglePAC performs competitively. For a label size of 2 and minimum dataset length of 22, JugglePAC achieves a latency of approximately 1.077 $\mu$s, comparable to or better than most previous designs such as DSA and SSA [13]. Its throughput is also high, outperforming designs like DB [19] and BTTP [20], especially with larger datasets.
JugglePAC operates at a frequency of 208 MHz, exceeding many previous designs like FPACC [16] and FCBT [13]. It achieves the lowest ”Slices ×$\times\mu$s” score, reflecting superior efficiency in balancing area and performance.
The full pipelining of JugglePAC enhances its performance and resource utilization, making it effective in high-speed operations. However, using a single floating-point adder might limit performance in cases requiring multiple adders. JugglePAC’s minimum dataset length varies with label size, ensuring accurate accumulation but potentially limiting flexibility. Future work could explore the incorporation of multiple adders to address these limitations and enhance performance.
Overall, JugglePAC represents a significant advancement in floating-point accumulation circuits, with strong performance across key metrics. The results highlight its potential for various applications and provide a basis for future research and optimization.
TABLE II: Comparison with previously proposed accumulation circuits.
Design Label Min. Dataset Adders Slices BRAMs Frequency Total Latency Slices×$\times\mu$s FPGA
Size Length (MHz) clock cycles $\mu$s
MFPA [14] - - 4 4,991 2 207 198 0.957 4,776 XC2VP30
AeMFPA [14] - - 2 3,130 14 204 198 0.970 3,036
$Ae^{2}$ [14] - - 2 3,737 2 144 198 1.370 5,120
FAAC [15] - - 3 6,252 0 199 176 1.086 6,790
FCBT [13] - - 2 2,859 10 170 $\leq$ 475 $\leq$ 2.794 7,988
DSA [13] - - 2 2,215 3 142 232 1.634 3,619
SSA [13] - - 1 1,804 6 165 $\leq$ 520 $\leq$ 3.152 5,686
DB [19] - - 1 1,749 6 188 $\leq$ 199 $\leq$ 1.058 1850
JugglePAC 1 74 1 1,439 0 208 $\leq$ 220 $\leq$ 1.058 1,522
JugglePAC 2 22 1 1,796 0 208 $\leq$ 224 $\leq$ 1.077 1,934
JugglePAC 3 10 1 2,343 0 208 $\leq$ 224 $\leq$ 1.077 2,523
FPACC [16] - - - 683 - 247 - - - VC5VSX50T
BTTP [20] - - 1 648 9.5 305 - - - XC5VLX110T
JugglePAC 2 22 1 578 0 334 $\leq$224 $\leq$ 0.671 388
V Conclusion
Accumulation is a fundamental operation, which appears in many types of computational workloads. In the case of high-throughput floating-point accumulation, complexities arise due to pipelining, especially when the system must handle consecutive datasets of varying lengths while producing results in the input order. Existing solutions often introduce significant overheads, while resulting in a reduced clock frequency or increased area complexity. In this work, we have introduced JugglePAC, a novel fully pipelined reduction circuit designed to overcome these challenges. JugglePAC leverages a single floating-point adder and yet efficiently manages accumulation tasks. Implemented and evaluated across multiple FPGAs, JugglePAC consistently delivers superior results in area and timing simultaneously, while previous works deliver either competitive area or timing, but not both at the same time.